# PCG-Informed Neural Solvers for High-Resolution Homogenization of Periodic Microstructures

### v1

@@ Abstract.
To address these challenges, we present CGiNS, a preconditioned-conjugate-gradient-solver-informed neural network for solving homogenization problems. CGiNS leverages sparse and periodic 3D convolution to enable high-resolution learning while ensuring structural periodicity. It features a multi-level network architecture that facilitates effective learning across different scales and employs minimum potential energy as label-free loss functions for self-supervised learning. The integrated preconditioned conjugate gradient iterations ensure that the network provides PCG-friendly initial solutions for fast convergence and high accuracy. Additionally, CGiNS imposes a global displacement constraint to ensure physical consistency, addressing a key limitation in prior methods that rely on Dirichlet anchors.
Evaluated on large-scale datasets with diverse topologies and material configurations, CGiNS achieves state-of-the-art accuracy (relative error below $1\%$) and outperforms both learning-based baselines and GPU-accelerated numerical solvers. Notably, it delivers 2×2$\times$2 × to 10×10$\times$10 × speedups over traditional methods while maintaining physically reliable predictions at resolutions up to $512^{3}$.
@@ 1. Introduction
Microstructures, prevalent in natural materials, have garnered significant interdisciplinary attention for their ability to achieve superior mechanical performance with minimal mass density, particularly in applications requiring high strength-to-weight and stiffness-to-weight ratios (Brandon and Kaplan, 2013). The advancements in additive manufacturing (AM) technologies have significantly improved microstructure fabrication, enabling the precise creation of geometrically complex architectures with customizable topologies and diverse base materials (Askari et al., 2020; Bajaj et al., 2020; Liu et al., 2022b; Tan et al., 2020). These developments have driven the widespread adoption of periodic microstructures in applications such as lightweight aerospace components and biomedical implants (Li, 1996).
@@
Homogenization theory provides a foundational framework for linking microscale material behavior to macroscopic effective properties in periodic media. At its core, homogenization reduces to solving a linear elasticity boundary value problem – i.e. , a system of partial differential equations (PDEs) that enforce mechanical equilibrium under periodic boundary conditions (Babuška, 1976; Terada et al., 2000; Bensoussan et al., 2011; Li, 1996; Torquato et al., 2002). By resolving these PDEs on representative unit cells, homogenization enables the derivation of continuum-scale constitutive laws without explicitly modeling the full microstructure.
Among available techniques, the finite element method (FEM) remains the most widely adopted approach due to its mesh flexibility and robustness across complex topologies (Omori and Kotera, 2007; Keyes et al., 2013; Solin, 2006). However, FEM-based homogenization typically involves the solution of large-scale sparse linear systems, where fine discretization required for accuracy imposes considerable computational and memory costs (Barkanov, 2001; Saeb et al., 2016). Iterative solvers, such as conjugate gradient (CG) methods and multigrid solvers, are often employed to alleviate computational costs through GPU acceleration, but they remain slow to converge, particularly for high-resolution discretizations. While Fast Fourier Transform (FFT)-based solvers are efficient for regular-grid periodic problems (Schneider, 2021; Chen et al., 2020), they suffer from convergence issues and numerical oscillations when applied to high-contrast microstructures (Schneider et al., 2016; Lucarini et al., 2022).
Figure 1. Overview of the CGiNS framework. This framework is illustrated using periodic lattice structures with a target resolution of $n=128$. Given a periodic microstructure, CGiNS voxelizes the input domain and applies soft-voxel padding to enhance boundary fidelity. The geometry and base material tensor are processed through a sparse convolutional encoder to extract multi-resolution features $\mathcal{Q}^{l}$. At each level $l$, a U-Net block refines these features, which are then decoded into displacement predictions $u^{l}$ and corrected via global translation to satisfy displacement constraints. A $t$-step preconditioned conjugate gradient (PCG) smoother eliminates high-frequency errors. Coarse-to-fine trilinear interpolation propagates feature information across levels, while energy-based losses $\mathcal{L}^{l}$ are computed at each resolution for self-supervised training. All computations operate on sparse tensors to ensure scalability to high-resolution structures.
@@
These approaches aim to fully replace traditional solvers by directly approximating PDE solutions using techniques such as physics-informed neural networks (PINNs) (Raissi et al., 2019; Cai et al., 2021a, b), neural operators (Fourier neural operator and DeepONet)(Li et al., 2020; Lu et al., 2021, 2019; Zappala et al., 2024; Cao et al., 2024), and emerging attempts to leverage large language models (LLMs) for PDE analysis (Zhou et al., 2024; Bhatnagar et al., 2025). While promising, these methods are primarily tested on relatively simple PDEs such as Poisson equations. Their applicability to more complex physical systems, such as homogenization PDEs with multi-scale coefficients and boundary coupling, remains underexplored. Furthermore, neural operator frameworks often require large-scale labeled datasets for training, and generating such data – especially for high-resolution or physics-specific problems – can be computationally prohibitive.
@@
These methods integrate neural networks with numerical solvers to enhance performance while retaining the robustness of traditional methods. Neural networks have been applied to accelerate intermediate solution steps (Chen et al., 2022; Han et al., 2024), predict efficient preconditioner (Trifonov et al., 2024; Azulay and Treister, 2022; Zhang et al., 2024), and improve initial guesses for iterative solvers (Zou et al., 2025; Huang et al., 2020; Pestourie et al., 2023). While the latter is conceptually related to our approach, the neural and numerical components in these methods remain loosely coupled – typically trained separately and connected in a sequential, non-differentiable manner.
@@
Existing neural network-based methods for microstructure homogenization, such as PH-Net (Peng et al., 2022) and Label-free (Zhu et al., 2024), directly approximate solutions to linear systems from geometric input. While effective on simple lattice structures, these models degrade significantly on complex topologies, often exhibiting instability and reduced accuracy (as shown in Section 2.2). Most results are reported on small datasets without large-scale evaluation, limiting generalization. Current models also struggle with high-resolution inputs due to memory and compute bottlenecks. Boundary conditions are typically imposed through weak supervision, such as loss penalties, rather than being explicitly encoded into the model architecture. In particular, Dirichlet conditions are often enforced by arbitrarily fixing a single node (anchor node) to zero displacement – an approach that may be sufficient for regular geometries but becomes unreliable for topologically complex microstructures, where the anchor point cannot be consistently or meaningfully inferred. This often leads to violation of physical constraints and further amplifies prediction error. We also found that the reported efficiency gains are often based on biased hardware comparisons and convergence conditions, similar to the observations by (McGreivy and Hakim, 2024). These deficiencies highlight the core challenges in neural microstructure homogenization: achieving robust high-accuracy predictions across diverse and complex structures, ensuring boundary condition consistency, scaling to high resolutions, and maintaining computational efficiency without sacrificing generalization.
@@
In this work, we present a novel conjugate-gradient-solver-informed neural framework – CGiNS– designed for the rapid and scalable simulation of homogenized linear elastic properties in periodic microstructures. Fig. 1 illustrates the overall architecture of CGiNS. Leveraging sparse 3D convolutions and periodic encoding, CGiNS predicts numerical solutions to homogenization PDEs directly from voxelized geometries and base material parameters, scaling up to resolutions of $512^{3}$.
@@
To ensure convergence and physical consistency, CGiNS embeds preconditioned conjugate gradient (PCG) iterations into the network architecture, enabling high-quality initializations that accelerate convergence while maintaining solver stability. The hybrid multi-level framework supports hierarchical feature extraction and resolution-aware interpolation, improving accuracy across structural scales. Additionally, a physics-informed variational loss based on minimum potential energy allows self-supervised training without labeled displacement fields. A global displacement constraint is also introduced to enforce boundary correctness without relying on heuristic anchors.
@@
To enable reproducible and rigorous evaluation, we construct three benchmark datasets featuring diverse geometric topologies and mechanical behaviors. Across all datasets, CGiNS achieves relative elasticity tensor errors below $1\%$, while outperforming both learning-based and GPU-accelerated numerical solvers by 2×2$\times$2 ×10×10$\times$10 × in runtime – demonstrating its superior accuracy, generalization, and computational efficiency.
@@ 2.1. Overview of CGiNS
Solving Eq. 2 is computationally expensive, particularly for high-resolution or topologically complex microstructures. To address this challenge, we propose CGiNS. (Further details on the homogenization setup and symbol definitions are provided in Supplementary.
@@
Fig. 1 illustrates the end-to-end architecture of CGiNS. The input microstructure is voxelized with soft voxel padding to enhance surface representation. Geometric and material properties are encoded via sparse convolutions into hierarchical features. A U-Net backbone transforms these features across resolutions using trilinear interpolation. At each level, displacement predictions are decoded, normalized via a global translation constraint, and refined via a $t$-step preconditioned conjugate gradient iterations (PCG module). Importantly, no ground-truth labels are required–CGiNS is trained using physically grounded energy-based loss at each resolution, ensuring stable convergence even for complex topologies. All modules are implemented in sparse format, enabling efficient training on resolutions up to $512^{3}$.
@@
We conduct all training and evaluation using three representative microstructure datasets (TPMS, PSL and Truss) with increasing geometric complexity. The TPMS dataset contains triply periodic minimal surface structures characterized by smooth, continuous morphologies and is widely used due to its regularity and predictable mechanics. In contrast, the PSL and Truss datasets exhibit complex topologies with sharp geometric variations and connectivity discontinuities, providing more realistic and challenging scenarios.
@@ 2.2. Accuracy and robustness of CGiNS
[…]
@@
2.3. Runtime efficiency compared to GPU-based solvers
To evaluate the computational efficiency of CGiNS, we compare it against two GPU-accelerated multigrid solvers: NVIDIA AmgX (Developer, 2017), which implements algebraic multigrid (AMG), and Homo3D (Zhang et al., 2023), which incorporates geometric multigrid (GMG) strategies with CUDA optimization. These solvers decompose computation into two phases: system matrix assembly (Init.) and iterative solution via preconditioned conjugate gradient (Cal.). AmgX supports resolutions up to $n=256$, while Homo3D extends to $n=512$.
@@ 2.3. Runtime efficiency compared to GPU-based solvers
Table 2. Performance of CGiNS on different structures and different materials. The volume fraction of the test models ranges from $10\%$ to $40\%$, with each subset containing 3000 microstructure models. $\|r\|$denotes the residual, divided by $10^{-3}$. For each group of test data, we record the maximum error $\delta_{Max}$, average error $\delta_{Mean}$, maximum residual $\|r\|_{Max}$, and average residual $\|r\|_{Mean}$.
@@
Table 3. Runtime and resolution scaling across solvers. We compare CGiNS with two numerical baselines: AmgX and Homo3D. All methods are evaluated on identical datasets with matched degrees of freedom (DOF), and residual convergence thresholds are aligned to those reached by CGiNS for fair comparison. ”Init.” denotes the time required to assemble system matrices or stencils (matrix assembly for AmgX, stencil preparation for Homo3D, and feature encoding for CGiNS). ”Cal.” refers to solver execution time: for AmgX and Homo3D, it measures iterative solving time; for CGiNS, it includes both neural inference and PCG iterations. At high resolutions ($n=512$), AmgX exceeds memory limits and is excluded.
Dataset Res. Avg. DOF AmgX (Developer, 2017) Homo3D (Zhang et al., 2023) CGiNS
Init. (s) Cal. Total Init. Cal. Total Init. Cal. Total Improve.
TPMS 64 $251K$ 0.04 0.27 0.32 0.62 0.25 0.87 0.05 0.09 0.14 2.13×2.13$\times$2.13 × / 5.89×5.89$\times$5.89 ×
128 $2M$ 0.30 3.10 3.40 0.67 1.04 1.71 0.28 0.55 0.82 4.11×4.$1$ / 2.06×2.06$\times$2.06 ×
256 $16M$ 1.27 7.84 9.10 0.76 6.71 7.46 0.02 3.30 3.33 2.73×2.73$\times$2.73 × / 2.24×2.24$\times$2.24 ×
512 $128M$ / / / 1.04 66.92 67.96 0.28 7.58 7.86 / 8.64×8.64$\times$8.64 ×
PSL 64 $275K$ 0.05 0.44 0.49 0.62 0.37 0.99 0.06 0.09 0.15 3.11×3.$1$ / 6.23×6.23$\times$6.23 ×
128 $2M$ 0.36 3.10 3.46 0.67 1.16 1.83 0.30 0.57 0.87 3.97×3.97$\times$3.97 × / 2.10×2.10$\times$2.10 ×
256 $17M$ 1.32 15.61 16.93 0.75 6.28 7.03 0.02 3.30 3.32 5.09×5.09$\times$5.09 × / 2.12×2.12$\times$2.12 ×
512 $136M$ / / / 1.04 80.25 81.29 0.34 10.88 11.22 / 7.24×7.24$\times$7.24 ×
Truss 64 $253K$ 0.05 0.41 0.46 0.62 0.39 1.01 0.06 0.09 0.15 2.95×2.95$\times$2.95 × / 6.38×6.38$\times$6.38 ×
128 $2M$ 0.30 3.22 3.52 0.67 1.17 1.84 0.26 0.59 0.85 4.11×4.$1$ / 2.15×2.15$\times$2.15 ×
256 $16M$ 1.37 19.32 20.69 0.75 7.18 7.93 0.02 3.30 3.32 6.22×6.$2$ / 2.38×2.38$\times$2.38 ×
512 $131M$ / / / 1.03 10.38 104.89 0.30 9.93 10.23 /10.25×10.25$\times$10.25 ×
@@
Figure 2. Accuracy–runtime and residual–runtime comparisons across resolutions. We compare CGiNS (red star) with two GPU-based solvers—AmgX (blue) and Homo3D (green)—on voxel grids of size $64^{3}$, $128^{3}$, $256^{3}$, and $512^{3}$. Top row: Relative error of the effective elasticity tensor $\delta(\%)$ versus wall-clock time, illustrating the accuracy–runtime tradeoff. Bottom row: Residual norm $\|r\|$ versus time, showing convergence behavior. All plots use logarithmic axes. At $512^{3}$, AmgX failed due to out-of-memory errors during matrix assembly and is therefore omitted.
As shown in Table 3, CGiNS achieves material property errors below $1\%$ across all tested resolutions, while offering 2×2$\times$2 × to 10×10$\times$10 × speedups relative to AmgX and Homo3D. Fig. 2 further reveals that although traditional solvers can continue reducing the residual norm $\|r\|$ through extended iterations, such improvements have negligible effect on the computed effective elasticity tensor $C^{H}$. This is because the mapping from displacement field $u$ to $C^{H}$ involves global energy integration and dot-product aggregation, which inherently smooth out local numerical perturbations and limit the practical benefit of over-solving the linear system.
This phenomenon has also been documented in the numerical homogenization literature (Schneider, 2021; Willot, 2015), where errors in $C^{H}$ below $1\%$ are widely accepted as sufficient for engineering applications. Together, these results confirm that CGiNS achieves not only substantial acceleration over classical solvers, but also converges to the necessary precision regime required for downstream material property prediction.

### v2

@@ Abstract.
To address these challenges, we present CGiNS, a geometry-aware preconditioned conjugate gradient solver informed neural network for solving homogenization problems. CGiNS leverages sparse and periodic 3D convolution to enable high-resolution learning while ensuring structural periodicity. It features a multi-level network architecture that facilitates effective learning across different scales and employs minimum potential energy as label-free loss functions for self-supervised learning. The integrated preconditioned conjugate gradient iterations ensure that the network provides PCG-friendly initial solutions for fast convergence and high accuracy. Additionally, CGiNS imposes a global displacement constraint to ensure physical consistency, addressing a key limitation in prior methods that rely on Dirichlet anchors.
Evaluated on large-scale datasets with diverse topologies and material configurations, CGiNS achieves state-of-the-art accuracy (relative error below $1\%$) and outperforms both learning-based baselines and GPU-accelerated numerical solvers. Notably, it delivers $10\times$ speedups over traditional methods while maintaining physically reliable predictions at $512^{3}$ resolution.
@@ 1. Introduction
Microstructures, prevalent in natural materials, have garnered significant interdisciplinary attention for their ability to achieve superior mechanical performance at minimal mass density, particularly in applications requiring high strength-to-weight and stiffness-to-weight ratios (Brandon and Kaplan, 2013). Periodic single-phase lattice microstructures constitute the dominant class implemented in additive-manufacturing (AM)–driven design and production: using a single certified feedstock simplifies fabrication and qualification, yields repeatable, designable effective properties, and scales reliably to large parts and databases of candidates (Askari et al., 2020; Bajaj et al., 2020; Liu et al., 2022b; Tan et al., 2020). As a result, single-phase periodic cells are the default choice in lightweighting, energy-absorption, and biomedical implant design (Li, 1996).
@@
Homogenization theory provides a foundational framework for linking microscale material behavior to macroscopic effective properties in periodic media. At its core, homogenization reduces to solving an elasticity boundary value problem – i.e. , a system of partial differential equations (PDEs) that enforce mechanical equilibrium under periodic boundary conditions (Babuška, 1976; Terada et al., 2000; Bensoussan et al., 2011; Li, 1996; Torquato et al., 2002). By resolving these PDEs on representative unit cells, homogenization enables the derivation of continuum-scale constitutive laws without explicitly modeling the full microstructure. In practice, multiple computational strategies have been developed. Fast Fourier transform (FFT)-based approaches are efficient for regular-grid periodic problems (Schneider, 2021; Chen et al., 2020), but they suffer from convergence issues and numerical oscillations when applied to high-contrast microstructures (Schneider et al., 2016; Lucarini et al., 2022). Theoretical homogenization methods, such as asymptotic expansion or mean-field theories (Sekkate et al., 2022), provide analytical approximations rather than fully numerical solutions. Among available techniques, the finite element method (FEM) remains the most widely adopted approach due to its mesh flexibility and robustness across complex topologies (Omori and Kotera, 2007; Keyes et al., 2013; Solin, 2006). However, FEM-based homogenization typically involves the solution of large-scale sparse linear systems, where fine discretization required for accuracy imposes considerable computational and memory costs (Barkanov, 2001; Saeb et al., 2016). Iterative solvers, such as conjugate gradient (CG) methods and multigrid solvers, are often employed to alleviate computational costs through GPU acceleration, but they remain slow to converge, particularly for high-resolution discretizations.
Figure 1. Overview of the CGiNS framework.This framework is illustrated using periodic lattice structures with a target resolution of $n$. Given a periodic microstructure, CGiNS voxelizes the input domain and applies soft-voxel padding to enhance boundary fidelity. The geometry and base material tensor are processed through a sparse convolutional encoder to extract multi-resolution features $\mathcal{Q}^{l}$. At each level $l$, a U-Net block refines these features, which are then decoded into displacement predictions $u^{l}$ and corrected via global translation to satisfy displacement constraints. A $t$-step preconditioned conjugate gradient (PCG) smoother eliminates high-frequency errors. Coarse-to-fine trilinear interpolation propagates feature information across levels, while energy-based losses $\mathcal{L}^{l}$ are computed at each resolution for self-supervised training. All computations operate on sparse tensors to ensure scalability to high-resolution structures.
@@
These approaches aim to fully replace traditional solvers by directly approximating PDE solutions using techniques such as physics-informed neural networks (PINNs) (Raissi et al., 2019; Cai et al., 2021a, b), neural operators (Fourier neural operator and DeepONet) (Li et al., 2020; Lu et al., 2021, 2019; Zappala et al., 2024; Cao et al., 2024), and emerging attempts to leverage large language models (LLMs) for PDE analysis (Zhou et al., 2024; Bhatnagar et al., 2025). While promising, these methods are primarily tested on relatively simple PDEs such as Poisson equations. Recent works have explored learning-based homogenization at the operator level (Bhattacharya et al., 2023, 2024), focusing on constructing reduced models of homogenized PDEs in idealized settings. In contrast, our work targets a solver-level challenge with industrial-grade accuracy: directly solving high-resolution 3D homogenization PDEs on voxelized microstructures under periodic boundary conditions, while embedding physical constraints and iterative solver consistency. Furthermore, neural operator frameworks often require large-scale labeled datasets for training, and generating such data – especially for high-resolution or physics-specific problems – can be computationally prohibitive.
@@
These methods integrate neural networks with numerical solvers to enhance performance while retaining the robustness of traditional methods. Neural networks have been applied to accelerate intermediate solution steps (Luz et al., 2020; Chen et al., 2022; Han et al., 2024), predict efficient preconditioner (Trifonov et al., 2024; Azulay and Treister, 2022; Zhang et al., 2024), and improve initial guesses for iterative solvers (Um et al., 2020; Zou et al., 2025; Huang et al., 2020; Pestourie et al., 2023). However, most of these approaches essentially operate from a matrix perspective, which inevitably discards underlying geometric information. We argue that the nature of this problem is a combination of geometry and numerics, and an effective solver should couple the two aspects more tightly. In our approach, the neural network component captures geometric characteristics and enforces problem-specific constraints such as periodicity, while the numerical part adopts PCG and iterative schemes. In this way, the network provides high-quality initial guesses to the numerical solver, and joint training further strengthens their coupling.
@@
Existing neural network-based methods for microstructure homogenization, such as PH-Net (Peng et al., 2022) and Label-free (Zhu et al., 2024), directly approximate solutions to linear systems from geometric inputs. While effective on simple lattice structures, these models degrade significantly on complex topologies, often exhibiting instability and reduced accuracy (as shown in Section 2.2). Most results are reported on small datasets without large-scale evaluation, limiting generalization. Current models also struggle with high-resolution inputs due to memory and compute bottlenecks. Boundary conditions are typically imposed through weak supervision, such as loss penalties, rather than being explicitly encoded into the model architecture. In particular, Dirichlet conditions are often enforced by arbitrarily fixing a single node (anchor node) to zero displacement – an approach that may be sufficient for regular geometries but becomes unreliable for topologically complex microstructures, where the anchor point cannot be consistently or meaningfully inferred. This often leads to violation of physical constraints and further amplifies prediction error. We also found that the reported efficiency gains are often based on biased hardware comparisons and convergence conditions, similar to the observations by (McGreivy and Hakim, 2024). These deficiencies highlight the core challenges in neural microstructure homogenization: achieving robust high-accuracy predictions across diverse and complex structures, ensuring boundary condition consistency, scaling to high resolutions, and maintaining computational efficiency without sacrificing generalization.
@@
In this work, we present a novel conjugate-gradient-solver-informed neural framework – CGiNS, designed for the rapid and scalable simulation of homogenized linear elastic properties in periodic microstructures. Fig. 1 illustrates the overall architecture of CGiNS. Leveraging sparse 3D convolutions and periodic encoding, CGiNS predicts numerical solutions to homogenization PDEs directly from voxelized geometries and base material parameters, scaling up to resolutions of $512^{3}$.
@@
To ensure convergence and physical consistency, CGiNS embeds preconditioned conjugate gradient (PCG) iterations into the network architecture, enabling high-quality initializations that accelerate convergence while maintaining solver stability. We use projection energy to quantify the alignment between the network initialization and the preconditioned CG system, providing key evidence for rapid convergence with only a few iterations. The hybrid multi-level framework supports hierarchical feature extraction and resolution-aware interpolation, improving accuracy across structural scales. Additionally, a physics-informed variational loss based on minimum potential energy enables self-supervised training without labeled displacement fields. A global displacement constraint is also introduced to enforce boundary correctness without relying on heuristic anchors.
@@
To enable reproducible and rigorous evaluation, we construct three benchmark datasets featuring diverse geometric topologies and mechanical behaviors. Across all datasets, CGiNS achieves relative elasticity tensor errors below $1\%$, while outperforming both learning-based and GPU-accelerated numerical solvers by $2\times$$10\times$ in runtime – demonstrating its superior accuracy, generalization, and computational efficiency.
@@ 2.1. Overview of CGiNS
Solving Eq. 2 is computationally expensive, particularly for high-resolution or topologically complex microstructures. To address this challenge, we propose CGiNS. (Further details on the homogenization setup and symbol definitions are provided in Supplementary.)
@@
Fig. 1 illustrates the architecture of CGiNS. The input microstructure is voxelized with soft voxel padding to enhance surface representation. Geometric and material properties are encoded via sparse convolutions into hierarchical features. A U-Net backbone transforms these features across resolutions using trilinear interpolation. At each level, displacement predictions are decoded, normalized via a global translation constraint, and refined via a $t$-step preconditioned conjugate gradient iterations (PCG module). Importantly, no ground-truth labels are required–CGiNS is trained using physically grounded energy-based loss at each resolution, ensuring stable convergence even for complex topologies. All modules are implemented in sparse format, enabling efficient training on resolutions up to $512^{3}$.
@@
We conduct all training and evaluation using three representative microstructure datasets (TPMS, PSL, and Truss) with increasing geometric complexity. The TPMS dataset contains triply periodic minimal surface structures characterized by smooth, continuous morphologies and is widely used due to its regularity and predictable mechanics. In contrast, the PSL and Truss datasets exhibit complex topologies with sharp geometric variations and connectivity discontinuities, providing more realistic and challenging scenarios.
@@ 2.2. Accuracy and robustness of CGiNS
Table 1. Performance of CGiNS on different structures. We compare CGiNS with SOTA learning-based approaches. To accommodate the resolution limit of all methods, we set $n=64$. The volume fraction of the test models ranges from $10\%$ to $40\%$, with each subset containing 3000 microstructure models. Since related works do not universally support different material configurations, both the test and training sets are configured with the same homogeneous material. $\delta$ denotes the relative error (in percentage, $\%$). $\|r\|$denotes the residual, divided by $10^{-3}$. For each test task, we report the maximum error $\delta_{Max}$, average error $\delta_{Mean}$, maximum residual $\|r\|_{Max}$, and average residual $\|r\|_{Mean}$.
@@
2.3. Runtime efficiency compared to linear equation solver
GPU-based numerical solver
To evaluate the computational efficiency of CGiNS, we compare it against two GPU-accelerated multigrid solvers: NVIDIA AmgX (Developer, 2017), which implements algebraic multigrid (AMG), and Homo3D (Zhang et al., 2023), which incorporates geometric multigrid (GMG) strategies with CUDA optimization. These solvers decompose computation into two phases: system matrix assembly (Init.) and iterative solution via preconditioned conjugate gradient (Cal.).
@@ 2.3. Runtime efficiency compared to linear equation solver
Table 2. Performance of CGiNS on different structures and different materials. The volume fraction of the test models ranges from $10\%$ to $40\%$, with each subset containing 3000 microstructure models. $\|r\|$denotes the residual, divided by $10^{3}$. For each group of test data, we record the maximum error $\delta_{Max}$, average error $\delta_{Mean}$, maximum residual $\|r\|_{Max}$, and average residual $\|r\|_{Mean}$.
@@
Table 3. Runtime and resolution scaling across solvers. We compare CGiNS with two purely numerical GPU multigrid solvers—AmgX (AMG) and Homo3D (GMG), and a numerical–neural hybrid method, GNN-AMG (Luz et al., 2020). All methods are evaluated on identical datasets with matched degrees of freedom (DOF), and the residual convergence thresholds are aligned to those reached by CGiNS for fair comparison. ”Init.” denotes the time required to assemble system matrices for AmgX, stencil preparation for Homo3D, graph feature building for GNN-AMG and feature encoding for CGiNS. Cal. refers to solver execution time: for AmgX and Homo3D, the time of their iterative solves; for GNN-AMG, the time of the AMG iterations; and for CGiNS, the time of neural inference plus PCG iterations. For GNN-AMG, S-rate is the success rate, i.e., the fraction of test instances that converge to the prescribed tolerance within the iteration budget.
Dataset Res. DOF AmgX (Developer, 2017) Homo3D (Zhang et al., 2023) GNN-AMG (Luz et al., 2020) CGiNS
Init. (s) Cal. Total Init. Cal. Total Init. Cal. Total S-rate Init. Cal. Total Improve.
TPMS 64 $251K$ 0.04 0.27 0.32 0.62 0.25 0.87 0.05 0.23 0.28 $87.11\%$ 0.05 0.09 0.14 $2.28\times$ / $6.21\times$ / $2.00\times$
128 $2M$ 0.30 3.10 3.40 0.67 1.04 1.71 0.28 2.90 3.18 $80.77\%$ 0.28 0.55 0.82 $4.11\times$ / $2.06\times$ / $3.87\times$
256 $16M$ 1.27 7.84 9.10 0.76 6.71 7.46 / / / / 0.02 3.30 3.33 $73\times / 2.2$$24\times$
512 $128M$ 2.28 80.52 82.80 1.04 66.92 67.96 / / / / 0.28 7.58 7.86 $10.24\times$ / $8.64\times$
PSL 64 $275K$ 0.05 0.44 0.49 0.62 0.37 0.99 0.06 1.07 1.13 $20.02\%$ 0.06 0.09 0.15 $3.11\times$ / $6.23\times$ / $7.53\times$
128 $2M$ 0.36 3.10 3.46 0.67 1.16 1.83 0.30 6.04 6.34 $10.48\%$ 0.30 0.57 0.87 $3.97\times$ / $2.10\times$ / $7.28\times$
256 $17M$ 1.32 15.61 16.93 0.75 6.28 7.03 / / / / 0.02 3.30 3.32 $5.09\times$ / $2.12\times$
512 $136M$ 2.34 120.50 122.84 1.04 80.25 81.29 / / / / 0.34 10.88 11.22 $10.77\times$ / $7.24\times$
Truss 64 $253K$ 0.05 0.41 0.46 0.62 0.39 1.01 0.06 1.30 1.36 $23.55\%$ 0.06 0.09 0.15 $2.95\times$ / $6.38\times$ / $9.07\times$
128 $2M$ 0.30 3.22 3.52 0.67 1.17 1.84 0.26 4.93 5.19 $17.03\%$ 0.26 0.59 0.85 $4.11\times$ / $2.15\times$ / $6.10\times$
256 $16M$ 1.37 19.32 20.69 0.75 7.18 7.93 / / / / 0.02 3.30 3.32 $6.22\times$ / $2.38\times$
512 $131M$ 2.30 121.49 123.47 1.03 103.86 104.89 / / / / 0.30 9.93 10.23 $11.85\times$ / $10.25\times$
@@
Figure 2. Accuracy–runtime and residual–runtime comparisons across resolutions. We compare CGiNS (red star) with two GPU-based solvers—AmgX (blue) and Homo3D (green)—on voxel grids of size $64^{3}$, $128^{3}$, $256^{3}$, and $512^{3}$. Top row: Relative error of the effective elasticity tensor $\delta(\%)$ versus wall-clock time, illustrating the accuracy–runtime tradeoff. Bottom row: Residual norm $\|r\|$ versus time, showing convergence behavior. All plots use logarithmic axes.
As shown in Table 3, CGiNS achieves material property errors below $1\%$ across all tested resolutions, while offering $2\times$ to $10\times$ speedups relative to AmgX and Homo3D. Fig. 2 further reveals that although traditional solvers can continue reducing the residual norm $\|r\|$ through extended iterations, such improvements have negligible effect on the computed effective elasticity tensor $C^{H}$. This is because the mapping from displacement field $u$ to $C^{H}$ involves global energy integration and dot-product aggregation, which inherently smooth out local numerical perturbations and limit the practical benefit of over-solving the linear system. This phenomenon has also been documented in the numerical homogenization literature (Schneider, 2021; Willot, 2015), where errors in $C^{H}$ below $1\%$ are widely accepted as sufficient for engineering applications. Together, these results confirm that CGiNS achieves not only substantial acceleration over classical solvers, but also converges to the necessary precision regime required for downstream material property prediction.
Numerical–neural hybrid solver
We also evaluate GNN-AMG (Luz et al., 2020), which delegates construction of the algebraic-multigrid prolongation operator $P$ to a message-passing GNN; as in the original design, we keep the AMG pipeline unchanged and replace only the values of $P$ with the network’s predictions. Because the method requires an explicit graph for $K$ and predicts all candidate entries of $P$, in our 3D periodic-lattice setting the maximum feasible resolution is limited to $128^{3}$ by GPU memory and runtime. A run is deemed successful if the linear solve reaches the prescribed tolerance.
As summarized in Table 3, on TPMS GNN-AMG achieves the highest success rate among our three datasets, and its solve time is lower than AmgX. In contrast, on PSL and Truss, the success rate is low and solve times exceed AmgX. These observations suggest that learning $P$ directly is more effective on simpler geometries, whereas for complex and highly variable geometries it often fails to converge.